# Anomalous dimensions from conformal field theory: generalized $\phi^{2n+1}$ theories

### v2

@@ 3.1 Multiplet recombination
Figure 4: An $\mathbb{R}^{3}$ projection of the $\mathbb{C}^{2}$ Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ at $N=3$. The surfaces are the Riemann surfaces of the functions $\ell(\Delta)$. The imaginary part of $\ell$ is shown by color. Red regions correspond to $\text{Im}[\ell]>0$, and blue regions are associated with $\text{Im}[\ell]<0$. The two turquoise points represent two solutions to $\ell=0$, which are $\Delta_{\pm}$ in (73). The plot is symmetric with respect to the plane $\text{Re}[\Delta-d/2]=0$. We have chosen the parameters $k=1$ and $\epsilon=0.3$.
Figure 5: An $\mathbb{R}^{3}$ projection of the $\mathbb{C}^{2}$ Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ at $N=3$. The peaks at $\text{Re}[\ell]=-2$ and $\text{Im}[\ell]=0$ correspond to the $1/(\ell+2)$ pole in the anomalous dimension. We indicate the imaginary part of $\Delta-d/2$ by color. The regions with $\text{Im}[\Delta-d/2]>0$ are in red, and the regions with $\text{Im}[\Delta-d/2]<0$ are in blue. Two of the solutions to $\Delta=d/2$ are $\ell_{\pm}$ in (74). They form a complex conjugate pair, and are indicated by the two turquoise points. We have used $k=1$ and $\epsilon=0.3$.
@@
Figure 7: An $\mathbb{R}^{3}$ projection of the $\mathbb{C}^{2}$ Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ at $N=6/5$. The peaks at $\text{Re}[\ell]=-2$, $\text{Im}[\ell]=0$ is due to the $1/(\ell+2)$ pole in the anomalous dimension. The imaginary part of $\ell$ is shown by color: red regions implies $\text{Im}[\ell]>0$, and blue regions indicate $\text{Im}[\ell]<0$. The turquoise points mark the solutions $\ell_{\pm}^{*}$ in (74). The plot is made at $k=1$ and $\epsilon=0.3$.
@@
$\ell^{*}_{\pm}$ $=-1\pm i\,\sqrt{\frac{2R^{\prime}}{3N-7}}\;\epsilon^{1/2}+O(\epsilon)\,,$ (74)
$\ell^{*}_{3}$ $=-2-\frac{2R^{\prime}}{3N-7}\,\epsilon+\frac{2R^{\prime}(12R^{\prime}+11N-23)}{3(3N-7)^{2}}\epsilon^{2}+O(\epsilon^{3})\,.$ (75)
@@
Let us discuss the case of the trivial representation, where $R^{\prime}=4(N-1)$. For $N>7/3$, the solutions $\ell^{*}_{\pm}$ correspond to complex spins, so they do not show up in the real $(\Delta-d/2,\ell)$ plane. For $31<N<7/$ 3, there are three intersection points, as in figure 3, since $R^{\prime}$ is negative and $\ell^{*}_{\pm}$ are real. At $N=1$, we have $R^{\prime}=0$ and $\gamma_{1}=O(\epsilon^{3/2})$, meaning that the anomalous dimensions are zero to order $\epsilon^{1}$. So the $N=1$ trajectory reduces to that in the free theory. [^22] Strictly speaking, the free trajectory should not include the horizontal line $\ell=-2$. The $N<1$ plots are similar to the $N>7/3$ ones, where $\ell^{*}_{\pm}$ are complex. Furthermore, we show the projections of $\mathbb{C}^{2}$ Chew-Frautschi plots in figures 4, 5, 6 and 7. For the standard representation, $R^{\prime}$ changes sign at $N=2$. There is one intersection point with the vertical axis for $N>7/3$ as well as for $N<2$, and there are three intersection points for $32<N<7/$ 3. For the remaining two representations, $R^{\prime}$ is always negative. There are only two scenarios: $N>7/3$ and $N<7/3$. The former gives the trajectories with three intersection points on the vertical axis, and the latter yields the trajectories with one intersection point.
@@
Now, we briefly comment on the $k>1$ Chew-Frautschi plots. One difference is the presence of more curves, which arise from the deformation of more horizontal lines in the Gaussian limit. Another difference is that we can consider trajectories with higher $m$. For each $m$, there are more solutions to $\Delta=d/2$, so the Regge trajectory can have more intersections with the vertical axis. Let us first discuss the Gaussian limit, where the scaling dimensions do not depend on the $S_{N+1}$ representations. For a positive integer $k$, the horizontal lines are given by $\frac{(1-3k-\ell-m)_{2k}}{\ell+k+2m}=0$, implying that there are $12k-$ 1 such lines. The diagonal lines are translations of those in the $k=1$ case. In the interacting theory, the trajectories are deformed into disconnected pieces. For example, the $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ trajectory in the $k=2$ theory and its Gaussian limit are shown in figure 8. As in the $k=1$ case, the Chew-Frautschi plots in the interacting theory depend on $N$. For both $m=0$ and $m=1$, there are three scenarios: $N>65/37$, $371<N<65/$ 37 and $N<1$. We present the Chew-Frautschi plots with different values of $N$ in figures 9, 10, 11 and 12. The Chew-Frautschi plots with $N<1$ are similar to case of $N>65/37$.
@@
Figure 8: Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ in the $k=2$ theory. The interacting case with $N=2$ is shown in red. In the Gaussian limit, there are two dashed diagonal lines given by $(\ell+2)^{2}=(\Delta-6)^{2}$, and there exist three dashed horizontal lines at $\ell=-3,-4,-5$.
@@
Figure 9: In the $k=2$ theory, the Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ are shown. These trajectories correspond to the case of $N>65/37$. We have set $\epsilon=10^{-3}$.
@@
Figure 10: Chew-Frautschi plot of the $k=2$ interacting theory. The Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ with $371<N<65/$ 37 are shown. We have used $\epsilon=10^{-3}$.
@@ 3.2 Analytic bootstrap
In this subsection, we use the analytic bootstrap to check the results from the multiplet recombination. As in subsection 2.2, we consider the crossing constraint on the four-point function
@@
Then, the crossing equations are written as (D), (D), (D) and (D). The functions $g^{(i)}(z,\bar{z})$ decompose into conformal blocks associated with the operators in the trivial, standard, symmetric traceless and antisymmetric representations. As in subsection 2.2, we focus on the leading order in the $z\rightarrow 0$ expansion, and we verify the multiplet-recombination results for the $m=0$ trajectory.
@@
$\tilde{\gamma}_{\mathcal{J}}\big{|}_{m=0}=\frac{P_{\phi}^{\text{Potts}}}{P_{\phi}^{\text{YL}}}\tilde{\gamma}_{\mathcal{J}}^{\text{YL}}\big{|}_{m=0}\times\begin{cases}\;\frac{(N-1)(N+1)^{2}}{N}\times N\qquad&\text{trivial}\\ \;\frac{N-2}{N-1}\times(N-1)(N+1)^{2}&\text{standard}\\ \;-(N+1)^{2}\times 1&\text{symmetric traceless}\\ \;(N+1)^{2}\times(-1)&\text{antisymmetric}\,,\end{cases}$ (79)
where $P_{\phi}^{\text{YL}}$ and $\tilde{\gamma}_{\mathcal{J}}^{\text{YL}}$ are the Yang-Lee data. This means that we consider the matching of the lightcone singularity in the Yang-Lee case, then we add the $N$-dependent factors and replace the Yang-Lee OPE coefficients with those in the Potts model. Using the OPE coefficients (24) and (63), [^24] According to (37), the OPE coefficients with different normalizations are related by $P_{\phi}^{\text{YL}}=\left(\lambda_{\phi\phi\phi}^{\text{YL}}\right)^{2}$ and $P_{\phi}=\lambda_{\phi\phi\phi}^{2}$. we verify that the result agrees with (69) from the multiplet recombination. We have also examined the $m=1$ trajectory in appendix B, and we confirm that the multiplet-recombination results are consistent with crossing symmetry.
@@ 4 Generalized $\phi^{2n+1}$ theory ($n>1$)
In this section, we study the $\phi^{2n+1}$ deformation of the free higher-derivative theory, with the integer $n>1$. To avoid derivative interactions, we require that $k$ and $12n-$ 1 have no common divisor. [^25] This is similar to the case of the $\phi^{2n}$ theory, where $k$ and $n-1$ have no common divisor Safari:2017irw ; Safari:2017tgs . The conditions for the $\phi^{2n+1}$ and $\phi^{2n}$ theories can be unified as follows. For a $\phi^{p}$ theory with integer $p$, we require that $k$ and $p-2$ have no common divisor other than 2. In this way, we can consistently consider the pure $\phi^{2n+1}$ interaction. The action at $d=d_{\text{u}}-\epsilon$ takes the form
$S=\int\mathrm{d}^{d}x\left(\phi\,\Box^{k}\phi+g\mu^{(2n-1)\epsilon/2}\phi^{2n+1}\right)\,,$ (80)
@@
This theory describes the generalized multicritical Yang-Lee edge singularity vonGehlen:1994rp ; Lencses:2022ira ; Lencses:2023evr ; Lencses:2024wib . For $k=1$, it was argued that the Lagrangian in (80) can be obtained from the Landau-Ginzburg Lagrangian of the multicritical Ising model. For the latter Lagrangian, a shift of the order parameter field by a purely imaginary constant generates $\mathbb{Z}_{2}$-odd terms. Then, one can tune the coupling constants to reach the multicritical point given by (80).262626 This procedure can also lead to a multicritical point with an interaction term consisting of an even number of $\phi$’s. See Lencses:2024wib for more details. This generalizes Fisher’s argument in Fisher:1978pf .
In this section, the CFT data is not completely fixed by the multiplet recombination. We determine the set of anomalous dimensions up to one unknown parameter $\alpha$, which can be fixed using the diagrammatic method in the canonical case. For arbitrary $\alpha$, We also check the consistency of the results with crossing symmetry using the analytic bootstrap.
@@ 4.1 Multiplet recombination
For $n>1$, the normalization $\alpha$ is not determined by the matching conditions generated from this equation, and $\alpha$ appears in the results for the anomalous dimensions, as well as the OPE coefficients. This distinction from the $n=1$ case is due to the vanishing of free correlator $\langle\phi^{2n}_{\text{f}}(x_{1})\phi_{\text{f}}(x_{2})\phi_{\text{f}}(x_{3})\rangle$ with $n>1$. In subsection 2.1, we see that the free theory data is an important input in the multiplet recombination approach. As there is less free theory input available here, the set of matching conditions is less constraining. Nevertheless, we can still obtain a consistent set of data using (82).
@@
The left-hand side has the same form as that of (7). The right-hand side is a free correlator which can be computed using Wick’s theorem. The result is
@@
$4^{2k}k!\left(\frac{d_{\text{u}}-2k}{2}\right)_{2k}(1-k)_{k-1}\lim_{\epsilon\rightarrow 0}\left(\alpha^{-2}\gamma_{\phi}\right)=(2n)!\,,$ (86)
where the upper critical dimension $d_{\text{u}}$ is given by (81). We obtain a relation between the anomalous dimension $\gamma_{\phi}$ and the unknown parameter $\alpha$:
$\gamma_{\phi}=\frac{(2n)!\,\alpha^{2}}{4^{2k}k!\left(\frac{2k}{2n-1}\right)_{2k}(1-k)_{k-1}}+O(\alpha^{3})\,.$ (87)
The next step is to consider three-point functions. Unfortunately, the matching condition for $\langle\Box^{k}\phi(x_{1})\phi(x_{2})\phi(x_{3})\rangle$ is not useful, as the corresponding free correlator $\langle\phi^{2n}_{\text{f}}(x_{1})\phi_{\text{f}}(x_{2})$ $\phi_{\text{f}}(x_{3})\rangle$ vanishes for $n>1$. Moreover, the matching condition for $\langle\Box^{k}\phi(x_{1})\Box^{k}\phi(x_{2})\Box^{k}\phi(x_{3})\rangle$ involves $\alpha$ and an unknown OPE coefficient $\lambda_{\phi\phi\phi}$. So we cannot determine $\alpha$ and $\gamma_{\phi}$ by following the procedure in section 2.1. Nevertheless, we cam still find the relation between $\lambda_{\phi\phi\phi}$ and $\alpha$ using the matching condition
@@
The left-hand side is computed in appendix A. The result takes the same form as the left-hand side of (2.1). We obtain the constraint on the OPE coefficient:
@@
The solution is
$\lambda_{\phi\phi\phi}=\frac{(-1)^{k}\left((2n)!\right)^{3}\alpha^{3}}{4^{3k}\left(n!\right)^{3}\left(\frac{k}{2n-1}\right)_{k}^{3}\left(-2k\frac{n-1}{2n-1}\right)_{3k}}+O(\alpha^{3})\,,$ (90)
which agrees with Codello:2017qek at $k=1$.
@@
$\gamma_{\mathcal{J}}=2\gamma_{\phi}-\frac{n(2n)!\,\alpha^{2}}{2^{4k-2}(2k-m-1)!(-m)_{m}\left(1-k\frac{2n+1}{2n-1}-\ell-m\right)_{2k}}+O(\alpha^{3})\,,$ (95)
where $\gamma_{\phi}$ is given in (87). For $k=1$, this result can also be derived from eq.(4.19) in Gliozzi:2017gzh . The special case (91) is associated with $\ell=m=0$. One can check that, to order $\alpha^{2}$, the anomalous dimension of the stress tensor $(m=k-1,\;\ell=2)$ vanishes, even though we do not know the explicit expression for $\alpha$. In other words, we cannot determine $\alpha$ by requiring that the stress tensor does not acquire anomalous dimension. [^27] We believe that the spinning recombination cannot fix $\alpha$ either. For example, at $k=1$, we find that $\lim_{\epsilon\rightarrow 0}(\alpha^{-2}\langle\Box\phi(x_{1})\phi(x_{2})\partial\mathcal{J}(x_{3}))\rangle$ is proportional to $\lim_{\epsilon\rightarrow 0}(\alpha^{-2}\gamma_{\mathcal{J}})$, so the matching condition should only yield a relation between $\gamma_{\mathcal{J}}$ and $\alpha$, and this relation ought to be the same as (95).
Since the anomalous dimensions are not fixed by our CFT analysis, we resort to inputs from the diagrammatic method. At $k=1$, Codello et al. computed the anomalous dimensions of some scalar operators in the $n>1$ theory Codello:2017epp . The result for $\gamma_{\phi}$ is expressed in terms of the critical coupling constant in their equation (5.7), which is consistent with our formula (87). Furthermore, they also obtained the critical coupling constant, so they fully determined the anomalous dimension $\gamma_{\phi}$. We use their results to fix $\alpha$  [^28] In Codello:2017epp , the general expression for the critical coupling constant $g$ in equation (5.6) may contain a typo, as it does not seem to reproduce their $n=2$ result in equation (3.2). Here we use their beta functionals to solve for $g$, and the result is consistent with their equation (3.2).
$\alpha^{2}=\frac{96\epsilon}{(2n-1)(2n)!\left(6-(2n+1)!(2n)!\,\Gamma\left(\frac{2n+1}{2n-1}\right)\mathcal{S}(n)\right)}+O(\epsilon^{2})\,,$ (96)

### v3

@@ 3.1 Multiplet recombination
Figure 3: Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ at $N=3$. The dashed lines indicate the $\epsilon\rightarrow 0$ limit of the Regge trajectory. We set $k=1$ and $\epsilon=10^{-3}$.
Figure 4: An $\mathbb{R}^{3}$ projection of the $\mathbb{C}^{2}$ Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ at $N=3$. The surfaces are the Riemann surfaces of the functions $\ell(\Delta)$. The imaginary part of $\ell$ is shown by color. Red regions correspond to $\text{Im}[\ell]>0$, and blue regions are associated with $\text{Im}[\ell]<0$. The two turquoise points represent two solutions to $\ell=0$, which are $\Delta_{\pm}$ in (73). The plot is symmetric with respect to the plane $\text{Re}[\Delta-d/2]=0$. We choose the parameters $k=1$ and $\epsilon=0.3$.
@@
Figure 6: An $\mathbb{R}^{3}$ projection of the $\mathbb{C}^{2}$ Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ at $N=3$. The peaks at $\text{Re}[\ell]=-2$ and $\text{Im}[\ell]=0$ correspond to the $1/(\ell+2)$ pole in the anomalous dimension. We indicate the imaginary part of $\Delta-d/2$ by color. The regions with $\text{Im}[\Delta-d/2]>0$ are in red, and the regions with $\text{Im}[\Delta-d/2]<0$ are in blue. Two of the solutions to $\Delta=d/2$ are $\ell_{\pm}$ in (74). They form a complex conjugate pair, and are indicated by the two turquoise points. We use $k=1$ and $\epsilon=0.3$.
@@
Figure 7: An $\mathbb{R}^{3}$ projection of the $\mathbb{C}^{2}$ Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ at $N=6/5$. The peaks at $\text{Re}[\ell]=-2$, $\text{Im}[\ell]=0$ is due to the $1/(\ell+2)$ pole in the anomalous dimension. The imaginary part of $\ell$ is shown by color: red region implies $\text{Im}[\ell]>0$, and blue regions indicate $\text{Im}[\ell]<0$. The turquoise points mark the solutions $\ell_{\pm}^{*}$ in (74). The plot is made at $k=1$ and $\epsilon=0.3$.
@@
$\ell^{*}_{\pm}$ $=-1\pm i\,\sqrt{\frac{2R^{\prime}}{3N-7}}\;\epsilon^{1/2}+\ldots\,,$ (74)
$\ell^{*}_{3}$ $=-2-\frac{2R^{\prime}}{3N-7}\,\epsilon+\frac{2R^{\prime}(12R^{\prime}+11N-23)}{3(3N-7)^{2}}\epsilon^{2}+\ldots\,.$ (75)
@@
Let us discuss the case of the trivial representation, where $R^{\prime}=4(N-1)$. For $N>7/3$, the solutions $\ell^{*}_{\pm}$ correspond to complex spins, so they do not show up in the real $(\Delta-d/2,\ell)$ plane. For $31<N<7/$ 3, there are three intersection points, as in figure 2, since $R^{\prime}$ is negative and $\ell^{*}_{\pm}$ are real. At $N=1$, we have $R^{\prime}=0$ and $\gamma_{\phi}=0\epsilon+\ldots$, so the anomalous dimensions of the bilinear operators are zero to order $\epsilon^{1}$. So the $N=1$ trajectory reduces to that in the free theory. [^30] Strictly speaking, the free trajectory should not include the horizontal line $\ell=-2$. The $N<1$ plots are similar to the $N>7/3$ ones, where $\ell^{*}_{\pm}$ are complex. Furthermore, we show the projections of $\mathbb{C}^{2}$ Chew-Frautschi plots in figures 4, 6, 5 and 7. For the standard representation, $R^{\prime}$ changes sign at $N=2$. There is one intersection point with the vertical axis for $N>7/3$ as well as for $N<2$, and there are three intersection points for $32<N<7/$ 3. For the remaining two representations, $R^{\prime}$ is always negative. There are only two scenarios: $N>7/3$ and $N<7/3$. The former gives the trajectories with three intersection points on the vertical axis, and the latter yields the trajectories with one intersection point.
@@
Now, we briefly comment on the $k>1$ Chew-Frautschi plots. One difference is the presence of more curves, which arise from the deformation of more horizontal lines in the Gaussian limit. Another difference is that we can consider trajectories with higher $m$. For each $m$, there are more solutions to $\Delta=d/2$, so the Regge trajectory can have more intersections with the vertical axis. Let us first discuss the Gaussian limit, where the scaling dimensions do not depend on the $S_{N+1}$ representations. For a positive integer $k$, the horizontal lines are given by $\frac{(k+\ell+m)_{2k}}{k+\ell+2m}=0$, implying that there are $12k-$ 1 such lines. The $45^{\circ}$ and $135^{\circ}$ lines are parallel to those in the $k=1$ case. In the interacting theory, the trajectories are deformed into disconnected pieces as the mixing problems are resolved. For example, the $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ trajectory in the $k=2$ theory and its Gaussian limit are shown in figure 8. As in the $k=1$ case, the Chew-Frautschi plots in the interacting theory depend on $N$. For both $m=0$ and $m=1$, there are three scenarios: $N>65/37$, $371<N<65/$ 37 and $N<1$. We present the Chew-Frautschi plots with different values of $N$ in figures 9, 10, 11, and 12. The Chew-Frautschi plots with $N<1$ are similar to the case of $N>65/37$.
@@
Figure 8: Chew-Frautschi plot for $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ in the $k=2$ theory. The interacting case with $N=2$ is shown in red. In the Gaussian limit, there are dashed $45^{\circ}$ and $135^{\circ}$ lines given by $(\ell+2)^{2}=(\Delta-6)^{2}$, and there exist three dashed horizontal lines at $\ell=-3,-4,-5$.
@@
Figure 9: In the $k=2$ theory, the Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ are shown. These trajectories correspond to the case of $N>65/37$. We set $\epsilon=10^{-3}$.
@@
Figure 10: Chew-Frautschi plot of the $k=2$ interacting theory. The Regge trajectories of $\mathcal{J}^{(m=0)}_{\ell,\,\text{trivial}}$ with $371<N<65/$ 37 are shown. We use $\epsilon=10^{-3}$.
@@ 3.2 Analytic bootstrap
In this subsection, we use the analytic bootstrap to check the results from the multiplet recombination. As in section 2.2, we consider the crossing constraint on the four-point function
@@
Then, the crossing equations are written as (D), (D), (D), and (D)., The functions $g^{(i)}(z,\bar{z})$ decompose into conformal blocks associated with the operators in the trivial, standard, symmetric traceless, and antisymmetric representations. As in section 2.2, we focus on the leading order in the $z\rightarrow 0$ expansion, and verify the multiplet-recombination results for the $m=0$ trajectory.
@@
$\tilde{\gamma}_{\mathcal{J}_{\ell}^{(0)}}=\frac{P_{\phi}^{\text{Potts}}}{P_{\phi}^{\text{YL}}}\tilde{\gamma}_{\mathcal{J}_{\ell}^{(0)}}^{\text{YL}}\times\begin{cases}\;(N-1)(N+1)^{2}\qquad&\text{trivial}\\ \;(N-2)(N+1)^{2}&\text{standard}\\ \;-(N+1)^{2}&\text{symmetric traceless}\\ \;-(N+1)^{2}&\text{antisymmetric}\,,\end{cases}$ (79)
where $P_{\phi}^{\text{YL}}$ and $\tilde{\gamma}_{\mathcal{J}_{\ell}^{(0)}}^{\text{YL}}$ are the Yang-Lee data. This means that we consider the matching of the lightcone singularity in the Yang-Lee case, and then we add the $N$-dependent factors and replace the Yang-Lee OPE coefficients with those in the Potts model. Using the OPE coefficients (22) and (61), [^32] According to (35), the OPE coefficients with different normalizations are related by $P_{\phi}^{\text{YL}}=\left(\lambda_{\phi\phi\phi}^{\text{YL}}\right)^{2}$ and $P_{\phi}=\lambda_{\phi\phi\phi}^{2}$. we obtain
$\tilde{\gamma}_{\mathcal{J}_{\ell}^{(0)}}=-\frac{R\,\alpha^{2}}{2^{4k-1}(2k-1)!(k+\ell)_{2k}}+\ldots\,,$ (80)
which is in agreement with the multiplet-recombination result (69) at $m=0$. We also examine the $m=1$ trajectory in appendix B, and confirm that the multiplet-recombination results are consistent with crossing symmetry. Similarly to (44), the agreement is valid for arbitrary $\alpha$.
@@ 4 Generalized $\phi^{2n+1}$ theory ($n>1$)
In this section, we study the $\phi^{2n+1}$ deformation of the free higher-derivative theory, where $n$ is larger than $1$. If $k$ and $12n-$ 1 have a common divisor, there may exist $\mathbb{Z}_{2}$-odd marginal derivative interactions, which can mix with $\phi^{2n+1}$. To avoid this, we assume that ($k,2n-1)$ is a pair of coprime integers.  [^33] This is similar to the case of the $\phi^{2n}$ theory, where $k$ and $n-1$ have no common divisor Safari:2017irw ; Safari:2017tgs . The conditions for the $\phi^{2n+1}$ and $\phi^{2n}$ theories can be unified as follows. For a $\phi^{p}$ theory with integer $p$, we require that $k$ and $p-2$ have no common divisor other than 2. In this way, we can consistently focus on the pure $\phi^{2n+1}$ interaction. The action at $d=d_{\text{u}}-\epsilon$ takes the form
$S\propto\int\mathrm{d}^{d}x\left(\phi\,\Box^{k}\phi+g\mu^{(2n-1)\epsilon/2}\phi^{2n+1}\right)\,,$ (81)
@@
This theory describes the generalized multicritical Yang-Lee edge singularity vonGehlen:1994rp ; Lencses:2022ira ; Lencses:2023evr ; Lencses:2024wib . For $k=1$, it was argued that the Lagrangian in (81) can be obtained from the Landau-Ginzburg Lagrangian of the multicritical Ising model. For the latter Lagrangian, a shift of the order parameter field by a purely imaginary constant generates $\mathbb{Z}_{2}$-odd terms. Then, one can tune the coupling constants to reach the multicritical point given by (81).343434 This procedure can also lead to a multicritical point with an interaction term consisting of an even number of $\phi$’s. See Lencses:2024wib for more details. This generalizes Fisher’s argument in Fisher:1978pf .
For $n>1$, the CFT data is not completely fixed by the multiplet recombination. We determine the set of anomalous dimensions up to one unknown parameter $\alpha$, which can be fixed using the diagrammatic method in the canonical case. For arbitrary $\alpha$, we also check the consistency of the results with crossing symmetry using the analytic bootstrap.
@@ 4.1 Multiplet recombination
For $n>1$, the normalization $\alpha$ is not determined by the matching conditions generated from this equation, and $\alpha$ appears in the results for the anomalous dimensions, as well as the OPE coefficients. This distinction from the $n=1$ case is due to the vanishing of the free-theory correlator $\langle\phi^{2n}_{\text{f}}(x_{1})\phi_{\text{f}}(x_{2})\phi_{\text{f}}(x_{3})\rangle$ for $n>1$. In the previous sections, we have seen that the free-theory data is an important input in the multiplet recombination approach. As there is less free-theory input available here, the set of matching conditions is less constraining. Nevertheless, we can still obtain a consistent set of data using (83).
@@
The left-hand side has the same form as that of (7). The right-hand side is a free correlator, which can be computed using Wick contractions. The result is
@@
$(-1)^{k-1}4^{2k}k!(k-1)!\left(\frac{d_{\text{u}}-2k}{2}\right)_{2k}\,\lim_{\epsilon\rightarrow 0}\left(\alpha^{-2}\gamma_{\phi}\right)=(2n)!\,,$ (87)
where the upper critical dimension $d_{\text{u}}$ is given by (82). This generalizes (13) in the $n=1$ case. We obtain a relation between the anomalous dimension $\gamma_{\phi}$ and the unknown parameter $\alpha$:
$\gamma_{\phi}=\frac{(-1)^{k-1}(2n)!\,\alpha^{2}}{4^{2k}k!(k-1)!\left(\frac{2k}{2n-1}\right)_{2k}}+\ldots\,,$ (88)
The next step is to consider three-point functions. Unfortunately, the matching condition for $\langle\Box^{k}\phi(x_{1})\phi(x_{2})\phi(x_{3})\rangle$ is not useful as the corresponding free correlator $\langle\phi^{2n}_{\text{f}}(x_{1})\phi_{\text{f}}(x_{2})$ $\phi_{\text{f}}(x_{3})\rangle$ vanishes for $n>1$. The matching condition for $\langle\Box^{k}\phi(x_{1})\Box^{k}\phi(x_{2})\Box^{k}\phi(x_{3})\rangle$ involves $\alpha$ and an unknown OPE coefficient $\lambda_{\phi\phi\phi}$, so we cannot determine $\alpha$ and $\gamma_{\phi}$ by following the procedure in section 2.1. Nevertheless, we can still deduce the relation between $\lambda_{\phi\phi\phi}$ and $\alpha$ using the matching condition
@@
Using (A), we obtain a constraint on the OPE coefficient:
@@
The equation above does not apply to $n=1$ because the factor $(\frac{d_{\text{u}}-6k}{4})_{3k}$ would vanish. As in (18), the correction to $\Delta_{\phi}$ should be taken into account for $n=1$, but the $n>1$ constraint (90) does not involve the anomalous dimension of $\phi$. The solution for (90) reads
$\lambda_{\phi\phi\phi}=\frac{(-1)^{k}\left((2n)!\right)^{3}\alpha^{3}}{4^{3k}\left(n!\right)^{3}\left(\frac{k}{2n-1}\right)_{k}^{3}\left(-2k\frac{n-1}{2n-1}\right)_{3k}}+\ldots\,,$ (91)
which agrees with the $k=1$ result in Codello:2017qek .
@@
$\gamma_{\mathcal{J}^{(m)}_{\ell}}=2\gamma_{\phi}-\frac{(-1)^{m}n^{2}(2n-1)!\,\alpha^{2}}{2^{4k-3}m!(2k-m-1)!\left(\ell+m-k\frac{2n-3}{2n-1}\right)_{2k}}+\ldots\,,$ (96)
where $\gamma_{\phi}$ is given in (88). This expression generalizes (27) in the case of $n=1$. For $k=1$, this result can also be derived from eq.(4.19) in Gliozzi:2017gzh . The special case (92) is associated with $\ell=m=0$. One can check that, to order $\alpha^{2}$, the anomalous dimension of the stress tensor $(m=k-1,\;\ell=2)$ vanishes, even though we do not know the explicit expression for $\alpha$. In other words, we cannot determine $\alpha$ by requiring that the stress tensor does not acquire an anomalous dimension. [^35] We believe that the spinning recombination cannot fix $\alpha$ either. For example, at $k=1$, we find that $\lim_{\epsilon\rightarrow 0}(\alpha^{-2}\langle\Box\phi(x_{1})\phi(x_{2})\partial\mathcal{J}(x_{3}))\rangle$ is proportional to $\lim_{\epsilon\rightarrow 0}(\alpha^{-2}\gamma_{\mathcal{J}})$, so the matching condition should only yield a relation between $\gamma_{\mathcal{J}}$ and $\alpha$, and this relation ought to be the same as (96).
Since the anomalous dimensions are not fixed by our CFT analysis, we resort to inputs from the diagrammatic method. At $k=1$, Codello et al. computed the anomalous dimensions of some scalar operators in the $n>1$ theory Codello:2017epp . The result for $\gamma_{\phi}$ is expressed in terms of the fixed-point coupling constant in their equation (5.7), which is consistent with our formula (88). Furthermore, they also obtained the fixed-point coupling constant, so they fully determined the anomalous dimension $\gamma_{\phi}$. We use their results to fix $\alpha$  [^36] In Codello:2017epp , the general expression for the fixed-point coupling constant $g$ in equation (5.6) may contain a typo, as it does not seem to reproduce their $n=2$ result in equation (3.2). Here we use their beta functionals to solve for $g$, and the result is consistent with their equation (3.2).
$\alpha^{2}=\frac{96\epsilon}{(2n-1)(2n)!\left(6-(2n+1)!(2n)!\,\Gamma(\frac{2n+1}{2n-1})\mathcal{S}(n)\right)}+\ldots\,,$ (97)